# DiffFuSR: Super-Resolution of all Sentinel-2 Multispectral Bands using Diffusion Models

### v1

@@ III-A1 Forward Diffusion Process
Following [30], the forward process is a Markov chain that progressively adds Gaussian noise to a ”clean” input image $\mathbf{x}_{0}$ over $T$ steps:
@@ III-A2 Reverse Denoising Process
The reverse denoising process is modeled as a Markov chain conditioned on the LR input $\mathbf{x}_{\text{LR}}$. This process iteratively denoises a sequence of latent variables $\mathbf{x}_{T},\dots,\mathbf{x}_{0}$ according to:
$p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T-1}|\mathbf{x}_{T},\mathbf{u},\mathbf{v})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{u},\mathbf{v}),$ (3)
@@
The variance $\sigma_{t}^{2}$ is typically pre-defined, while the mean $\boldsymbol{\mu}_{\boldsymbol{\theta}}$ is predicted by a denoising network, commonly a U-Net architecture [31], conditioned on the noisy latent $\mathbf{x}_{t}$, the timestep $t$, and the fixed context $\mathbf{u},\mathbf{v}$ The conditioning inputs $\mathbf{u}$ and $\mathbf{v}$ are fixed throughout the reverse process and derived from the observed LR image $\mathbf{x}_{\text{LR}}$ as follows:
@@
  • The global degradation embedding $\mathbf{v}=g_{\boldsymbol{\psi}}(\mathbf{x}_{\text{LR}})$ summarizes the underlying blur, noise, and downsampling characteristics of $\mathbf{x}_{\text{LR}}$. This embedding is computed using a separate encoder $g_{\boldsymbol{\psi}}$, trained via contrastive learning [32], and serves to modulate the denoising network through degradation-aware convolutions. The degradation embedding is injected at multiple positions in the network since it modulates the kernels of degradation-aware convolution layers.
The integration of $\mathbf{u}$ and $\mathbf{v}$ allows the model to incorporate both spatial structural details and degradation-specific behavior, thereby enabling accurate and blind SR.
@@ III-A3 Network Architecture
The proposed diffusion model is comprised of three main components (Fig. 2): (i) a denoising U-Net that predicts the clean high-resolution image from noisy input, (ii) a spatial encoder that extracts contextual features from the low-resolution image, and (iii) a degradation encoder trained with contrastive learning to summarize the degradation characteristics. We follow the design by Wu et al. [14].
@@
The core network $h_{\boldsymbol{\theta}}$ is a conditional U-Net that predicts the clean image $\mathbf{x}_{0}$ given a noisy input $\mathbf{x}_{t}$, timestep $t$, and conditioning vectors $\mathbf{u},\mathbf{v}$. To improve computational efficiency without losing information, we apply a pixel folding operation to $\mathbf{x}_{t}$ at the input [14]. This reduces spatial resolution by a factor $r$ while increasing the number of channels by $r^{2}$. Formally, this maps an image of dimension $C\times H\times W$ to $(C\cdot r^{2})\times(H/r)\times(W/r)$. The inverse of this operation is pixel shuffle as proposed in [33]. We use pixel shuffle to retrieve the full-sized final image. The folded input is passed through the U-Net, which consists of:
  • a downsampling path with multiple feature extraction blocks,
  • an upsampling path with nearest-neighbour interpolation and convolution, (i.e., spatial resolution is increased by replicating neighboring pixels, followed by refinement with convolution layers)
  • residual blocks that incorporate conditioning via timestep embedding and degradation information.
After denoising, the output is unfolded back to the original spatial resolution using the pixel shuffle operation.
@@
and then passed through a multi-layer perceptron (MLP) to produce the embedding $\mathbf{t}_{e}$. This embedding is added to intermediate feature maps throughout the U-Net [14].
@@
The LR input $\mathbf{x}_{\text{LR}}$ is processed by a spatial encoder $f_{\boldsymbol{\phi}}$, implemented as an RRDB-Net [26] consisting of 23 Residual-in-Residual Dense Blocks (RRDBs) with 64 feature channels. This architecture is known for its strong performance in SR tasks, particularly its ability to extract hierarchical spatial features without normalization layers. The encoder outputs a spatial feature map $\mathbf{u}=f_{\boldsymbol{\phi}}(\mathbf{x}_{\text{LR}})$, which serves as a global spatial conditioning signal for the denoising model.
Rather than being fused at multiple layers, the spatial feature map $\mathbf{u}$ is concatenated with the input of the U-Net denoiser at the first layer only. This provides the network with high-level spatial context from the LR image while keeping the fusion interface lightweight. During training, the decoder (a light weight CNN, not shown in Fig. 2) is attached to the encoder to predict a coarse high-resolution reconstruction $\hat{\mathbf{x}}_{0}$, supervised with an $\ell_{1}$ consistency loss. The decoder is discarded during inference.
@@
To model the implicit degradation of the LR image, we use a lightweight CNN $g_{\boldsymbol{\psi}}$ that maps an LR image patch $\mathbf{x}_{\text{LR}}$ to a global degradation vector $\mathbf{v}=g_{\boldsymbol{\psi}}(\mathbf{x}_{\text{LR}})$, following the contrastive learning formulation of [34]. The network consists of several convolutional layers, followed by global average pooling and a projection head. It is trained with a contrastive objective. In the contrastive setting, the goal is to train a model such that given two crops $\mathbf{x}_{q}$, $\mathbf{x}_{k}$ from the same LR image (i.e., with the same degradation), their embeddings $\mathbf{v}_{q}$, $\mathbf{v}_{k}$ should be similar, while embeddings from different images (assumed to have different degradations) should be dissimilar. Under this setting, the model learns about the degradation in an unsupervised manner.
The learned degradation vector $\mathbf{v}$ is injected into each residual block in the U-Net using a degradation-aware convolution (DAConv) layer, as proposed by Wu et al. [14], which adapts the convolutional filters based on the conditioning vector. In this approach, the convolutional filters are dynamically modulated based on the conditioning vector, allowing the network to adapt to different degradation types during inference.
@@ III-A4 Training Objective
  • Denoising loss: The diffusion model is trained to predict the clean high-resolution image $\mathbf{x}_{0}$ from its noisy version $\mathbf{x}_{t}$, conditioned on spatial and degradation features using the ELBO loss [30]:
@@
  • Contrastive loss: We use the InfoNCE loss for training. Given two patches $\mathbf{x}_{\text{LR}}$ and $\mathbf{x}_{\text{LR}}^{+}$ from the same image (positive pair), and a set of negatives $\{\mathbf{x}_{\text{LR}}^{-}\}$ from different images in the batch, the contrastive loss is defined as:
@@ III-A5 Sampling
Both the spatial encoder $f_{\boldsymbol{\phi}}$ and degradation encoder $g_{\boldsymbol{\psi}}$ receive the same degraded image as input. The randomized degradation ensures that the degradation vector $\mathbf{v}$ is learned from data without relying on ground truth kernels which is key component of the model’s blind SR capability. Next, we will describe how we degrade and prepare the HR images for training the diffusion model.
@@ III-A6 Simulating Sentinel-2 images
Due to the absence of HR Sentinel-2 imagery suitable for supervised learning, we use HR NAIP aerial images as a proxy to construct synthetic training pairs.
Previous works such as the SEN2NAIP study [3] leverage real, co-registered NAIP–Sentinel-2 pairs to train a harmonization model, comparing a deterministic neural network-based approach with a statistical method grounded in power-law transformations. While the authors did not claim a definitive advantage for either method, they noted that the neural network approach produced slightly sharper visual results for SR. In contrast, we found that learned harmonization does not consistently yield sharper outputs and introduces unnecessary computational overhead. Therefore, we adopt a simpler, statistics-based harmonization model using per-channel power-law transformations.
We generate synthetic LR Sentinel-2-like images by applying a two-stage transformation to NAIP data: (i) spectral harmonization to match Sentinel-2 reflectance distributions, and (ii) randomized degradation to simulate optical and sensor effects such as blur and noise. This strategy enables fully self-supervised training and supports blind SR without reliance on paired ground truth.
@@
Harmonization adjusts the spectral distribution of NAIP imagery to resemble that of Sentinel-2 RGB bands. This process mitigates radiometric discrepancies between the two sensors, enabling more effective cross-domain learning. We perform per-channel gamma correction using a power-law transformation:
@@
where $\mathbf{x}^{c}_{NP}$ denotes band $c$ of the NAIP image, $k=255$ denotes the dynamic range, and $\gamma_{c}$ is the channel-specific gamma value. These gamma values are not arbitrarily chosen; rather, they are derived empirically by aligning NAIP reflectance distributions to co-registered Sentinel-2 imagery. Following the methodology of Aybar et al. [3], optimal gamma values are computed for each channel across a large corpus of image pairs. A multivariate normal model is then fit to the resulting gamma vectors, and the median values are used to define a global correction profile. This ensures consistent spectral alignment during inference without the need for image-specific tuning.
@@
To emulate optical blur, we use both isotropic and anisotropic Gaussian kernels, with parameters sampled randomly for each training image. We select a fixed kernel size of $21\times 21$. During training, we use an anisotropic Gaussian blur type, whereas for validation, we use an isotropic Gaussian. We sample the smoothing parameter $\sigma$ from the range $\sigma\in[2.0,4.0]$, the eigenvalue range of the anisotropic Gaussian kernel as $\lambda_{1},\lambda_{2}\in[0.2,4.0]$, and the rotation angle as $\theta\in[0,\pi]$. Our strategy is different from Aybar et al. [3] since they suggested applying a fixed smoothing parameter ($\sigma$) for the blur kernel. From empirical evaluations, we concluded that a blind approach led to better results. In addition, we also found that using both isotropic and anisotropic blur is important.
@@ III-B Multispectral Band Fusion
Figure 3: Architecture of the multispectral fusion pipeline. At test time, each resolution group of 10 m, 20 m, and 60 m bands is processed by a dedicated fusion module that receives the native-resolution multispectral inputs along with the RGB image super-resolved to 2.5 m. These modules independently reconstruct each group at 2.5 m GSD. The outputs are then merged to produce a consistent 12-band super-resolved Sentinel-2 image. This process is applied during both training (using synthetic degraded inputs) and inference (using native-resolution bands), with shared architecture and supervision strategy across all groups. The scaling factor for degraded input for training is computed using the Wald protocol. For example, for the 60 m group, we aim to learn the mapping from 60 m to 2.5 m given a 2.5 m SR guidance signal. This corresponds to a 24× scale factor. During training, all bands (10 m, 20 m, and 60 m) are degraded by this factor, producing input resolutions of 240 m, 480 m, and 1440 m, respectively. At training time, the network reconstructs the degraded signal to 60 m GSD, supervised by the corresponding 60 m ground-truth reference.
To produce a consistent 2.5 m resolution across all Sentinel-2 bands, we design a learnable fusion approach that reconstructs HR multispectral bands by injecting spatial detail from the super-resolved RGB image (Fig. 3). This is formulated as a guided upsampling problem increasing resolution by pixel replication and feature refinement, where the RGB image provides high-frequency structural cues and the lower-resolution multispectral bands contribute the spectral content.
@@
such that the reconstructed data $\widehat{\mathbf{X}}_{\text{HR}}$ captures fine structural detail aligned with the super-resolved Sentinel-2 RGB image at 2.5 m GSD from the diffusion module, $\mathbf{x}_{0}$, while remaining spectrally consistent with the original $\mathbf{X}_{\text{LR}}$ data. Here $\boldsymbol{\nu}$ denotes the parameters of the fusion model. In practice, the fusion pipeline contains three different modules, one for each band group (10m, 20m, 60m), which are trained and tested simultaneously but completely independently of each other, as shown in Fig. 3.
@@ III-B1 Training Strategy and Degradation Simulation
The fusion model is trained in a fully self-supervised manner using only native Sentinel-2 data. No HR targets are assumed. Instead, we adopt a synthetic training strategy: the input bands are downsampled further from their native resolution, and the original band serves as the supervision target. This follows the Wald protocol in spirit, though adapted for groups with no native HR ground truth (e.g., 60 m bands).
To simulate realistic LR observations during training, each input band undergoes a degradation process comprising:
  • Blur: An optional isotropic Gaussian blur is applied to emulate sensor response.
  • Downsampling: Spatial resolution is reduced using a boxcar filter, implemented as average pooling with stride corresponding to the target scale. This approximates spatial averaging seen in optical sensor acquisition.
At test time, the fusion model receives inputs at the original resolution (10 m, 20 m, or 60 m), along with the super-resolved RGB image in 2.5 m resolution from the diffusion model. The super-resolved RGB input is not used during training. However, the network is expected to learn to generalize zero-shot to enhanced spatial priors from the super-resolved RGB image at 2.5 m [8].
@@ III-B2 Network Implementation
The fusion model $F_{\boldsymbol{\nu}}(\cdot,\cdot)$ is instantiated via a separate model for each resolution group (10 m, 20 m, and 60 m), each implemented as a two-branch convolutional network. One branch processes the degraded multispectral inputs, while the other processes a fusion signal derived from auxiliary inputs (e.g., RGB and NIR bands). These branches are merged and refined via a series of residual blocks with dense skip connections.
The inputs to each branch are preprocessed with bilinear interpolation to restore native resolution, after applying blur and boxcar degradation in the training phase. At inference time, the RGB input is substituted with the 2.5 m super-resolved output, allowing the model to exploit high-frequency guidance it has never seen during training.
The model is trained to minimize the $\ell_{1}$ reconstruction loss:
@@
$\mathcal{L}_{\text{fusion}}=\left\|\widehat{\mathbf{X}}_{\text{HR}}-\mathbf{X}_{\text{HR}}\right\|_{1},$
where $\mathbf{X}_{\text{HR}}$ is the native-resolution Sentinel-2 band used as the supervision target, and $\widehat{\mathbf{X}}_{\text{HR}}$ is the predicted fused output. This objective encourages the model to learn structure-preserving spectral fusion in a purely self-supervised fashion.

### v2

@@ III-A1 Forward Diffusion Process
Following [ho2020denoising], the forward process is a Markov chain that progressively adds Gaussian noise to a ”clean” input image $\mathbf{x}_{0}$ over $T$ steps:
@@ III-A2 Reverse Denoising Process
The reverse denoising process is modeled as a Markov chain conditioned on the LR input $\mathbf{x}_{\text{LR}}$. Starting from Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, this process iteratively denoises a sequence of latent variables $\mathbf{x}_{T},\dots,\mathbf{x}_{0}$ according to:
$p_{\boldsymbol{\theta}}(\mathbf{x}_{0:T}|\mathbf{u},\mathbf{v})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{u},\mathbf{v}),$ (3)
@@
The variance $\sigma_{t}^{2}$ is typically pre-defined, while the mean $\boldsymbol{\mu}_{\boldsymbol{\theta}}$ is predicted by a denoising network, commonly a U-Net architecture [ronneberger2015u], conditioned on the noisy latent $\mathbf{x}_{t}$, the timestep $t$, and the fixed context $\mathbf{u},\mathbf{v}$. The integration of $\mathbf{u}$ and $\mathbf{v}$ allows the model to incorporate both spatial structural details and degradation-specific behavior, thereby enabling accurate and blind SR.
The conditioning inputs $\mathbf{u}$ and $\mathbf{v}$ are fixed throughout the reverse process and derived from the observed LR image $\mathbf{x}_{\text{LR}}$ as follows:
@@
  • The global degradation embedding $\mathbf{v}=g_{\boldsymbol{\psi}}(\mathbf{x}_{\text{LR}})$ summarizes the underlying blur, noise, and downsampling characteristics of $\mathbf{x}_{\text{LR}}$. This embedding is computed using a separate encoder $g_{\boldsymbol{\psi}}$, trained via contrastive learning [he2020momentum].
@@ III-A3 Network Architecture
The proposed diffusion model is comprised of three main components (Fig. 2): (i) a denoising U-Net that predicts the clean HR image from noisy input, (ii) a spatial encoder that extracts contextual features from the low-resolution image, and (iii) a degradation encoder trained with contrastive learning to summarize the degradation characteristics. We follow the design by Wu et al. [wu2023conditional]:
@@
The core network $h_{\boldsymbol{\theta}}$ is a conditional U-Net that predicts the clean image $\mathbf{x}_{0}$ given a noisy input $\mathbf{x}_{t}$, timestep $t$, and conditioning vectors $\mathbf{u},\mathbf{v}$. To improve computational efficiency without losing information, we apply a pixel folding operation to $\mathbf{x}_{t}$ at the input [wu2023conditional]. This reduces spatial resolution by a factor $r$ while increasing the number of channels by $r^{2}$. Formally, this maps an image of dimension $C\times H\times W$ to $(C\cdot r^{2})\times(H/r)\times(W/r)$. The folded input is passed through the denoising U-Net [wu2023conditional]. After denoising, the output is unfolded back to the original spatial resolution using the pixel shuffle operation [shi2016real].
@@
and then passed through a multi-layer perceptron (MLP) to produce the embedding $\mathbf{t}_{e}$. This embedding is added to intermediate feature maps throughout the U-Net [wu2023conditional].
@@
The LR input $\mathbf{x}_{\text{LR}}$ is processed by a spatial encoder $f_{\boldsymbol{\phi}}$, implemented as an RRDB-Net consisting of 23 Residual-in-Residual Dense Blocks (RRDBs) with 64 feature channels [wang2018esrgan]. The encoder outputs a spatial feature map $\mathbf{u}=f_{\boldsymbol{\phi}}(\mathbf{x}_{\text{LR}})$, which serves as a global spatial conditioning signal for the denoising model.
Rather than being fused at multiple layers, the spatial feature map $\mathbf{u}$ is concatenated with the input of the U-Net denoiser at the first layer only. This provides the network with high-level spatial context from the LR image while keeping the fusion interface lightweight. During training, the decoder (a light weight CNN, not shown in Fig. 2) is attached to the encoder to predict a coarse HR reconstruction $\widehat{\mathbf{x}}_{0}$, supervised with an $\ell_{1}$ consistency loss. The decoder is discarded during inference.
@@
To model the implicit degradation of the LR image, we use a lightweight CNN $g_{\boldsymbol{\psi}}$ that maps an LR image patch $\mathbf{x}_{\text{LR}}$ to a global degradation vector $\mathbf{v}=g_{\boldsymbol{\psi}}(\mathbf{x}_{\text{LR}})$. The network consists of several convolutional layers, followed by global average pooling and a projection head. It is trained with a contrastive objective [wang2021unsupervised], where the goal is to train a model such that given two crops $\mathbf{x}_{q}$, $\mathbf{x}_{k}$ from the same LR image (i.e., with the same degradation), their embeddings $\mathbf{v}_{q}$, $\mathbf{v}_{k}$ should be similar, while embeddings from different images (assumed to have different degradations) should be dissimilar. Under this setting, the model learns about the degradation in an unsupervised manner.
The learned degradation vector $\mathbf{v}$ is injected into each residual block in the U-Net using a degradation-aware convolution (DAConv) layer [wu2023conditional]. In this approach, the convolutional filters are dynamically modulated based on the conditioning vector, allowing the network to adapt to different degradation types during inference.
@@ III-A4 Training Objective
  • Denoising loss: The diffusion model is trained to predict the clean HR image $\mathbf{x}_{0}$ from its noisy version $\mathbf{x}_{t}$, conditioned on spatial and degradation features using the ELBO loss [ho2020denoising]:
@@
  • Contrastive loss: We use the InfoNCE loss for training [wu2023conditional]. Given two patches $\mathbf{x}_{\text{LR}}$ and $\mathbf{x}_{\text{LR}}^{+}$ from the same image (positive pair), and a set of negatives $\{\mathbf{x}_{\text{LR}}^{-}\}$ from different images in the batch, the contrastive loss is defined as:
@@ III-A6 Simulating Sentinel-2 images
We adopt a simple, statistics-based harmonization model using per-channel power-law transformations to generate synthetic Sentinel-2-like images from NAIP images. This is achieved by a two-stage transformation: (i) spectral harmonization to match Sentinel-2 reflectance distributions, and (ii) randomized degradation to simulate optical and sensor effects such as blur and noise. This strategy enables fully self-supervised training and supports blind SR without reliance on paired ground truth.
@@
Harmonization adjusts the spectral distribution of NAIP imagery to resemble the Sentinel-2 RGB bands, enabling more effective cross-domain learning. We perform per-channel gamma correction using a power-law transformation:
@@
where $\mathbf{x}^{c}_{NP}$ denotes band $c$ of the NAIP image, $k=255$ denotes the dynamic range, and $\gamma_{c}$ is the channel-specific gamma value. These gamma values are derived empirically by aligning NAIP reflectance distributions to co-registered Sentinel-2 imagery. Optimal gamma values are computed for each channel across a large corpus of image pairs [aybar2024sen2naip]. A multivariate normal model is then fit to the resulting gamma vectors, and the median values are used to define a global correction profile. This ensures consistent spectral alignment during inference without the need for image-specific tuning.
@@
To emulate optical blur, we use both isotropic and anisotropic Gaussian kernels, with parameters sampled randomly for each training image. We select a fixed kernel size of $21\times 21$. During training, we use an anisotropic Gaussian blur type, whereas for validation, we use an isotropic Gaussian. We sample the smoothing parameter $\sigma$ from the range $\sigma\in[2.0,4.0]$, the eigenvalue range of the anisotropic Gaussian kernel as $\lambda_{1},\lambda_{2}\in[0.2,4.0]$, and the rotation angle as $\theta\in[0,\pi]$.
@@ III-B Multispectral Band Fusion
Figure 3: Architecture of the multispectral fusion pipeline. At test time, each resolution group of 10 m, 20 m, and 60 m bands is processed by a dedicated fusion module that receives the native-resolution multispectral inputs along with the RGB image super-resolved to 2.5 m. These modules independently reconstruct each group at 2.5 m GSD. The outputs are then merged to produce a consistent 12-band super-resolved Sentinel-2 image. This process is applied during both training (using synthetic degraded inputs) and inference (using native-resolution bands), with shared architecture and supervision strategy across all groups. The scaling factor for degraded input for training is computed using the Wald protocol. For example, for the 60 m group, we aim to learn the mapping from 60 m to 2.5 m given a 2.5 m SR guidance signal. This corresponds to a 24× scale factor. During training, 60m bands are degraded by this factor, producing input resolutions of 1440 m. At training time, the network reconstructs the degraded signal to 60 m GSD, supervised by the corresponding 60 m ground-truth reference.
Figure 4: Architecture of the proposed GLP-inspired [vivone2018full] fusion module. Each model variant for the 10 m, 20 m, and 60 m branches contains approximately 202k trainable parameters. The symbols denote basic operations: “+” indicates element-wise addition, “-” indicates subtraction, $\times$” indicates element-wise multiplication, and “c” indicates channel-wise concatenation.
To produce a consistent 2.5 m resolution across all Sentinel-2 bands, we design a learnable fusion approach that reconstructs HR multispectral bands by injecting spatial details from the super-resolved RGB image, $\mathbf{x}_{0}$ (Fig. 3). This is formulated as a guided upsampling problem, where the RGB image provides high-frequency structural cues and the LR multispectral bands contribute the spectral content.
@@
where $\widehat{\mathbf{X}}_{\text{HR}}$ is the reconstructed data at 2.5m GSD, and $\boldsymbol{\nu}$ denotes the parameters of the fusion model. In practice, the fusion pipeline contains three different modules, one for each band group (10m, 20m, 60m), which are trained and tested simultaneously but completely independently of each other (Fig. 3).
The three proposed fusion pipelines are based on the Generalized Laplacian Pyramid (GLP) with modulation transfer function (MTF) filters and a regression injection model (MTF-GLP) [vivone2018full, aiazzi2002context], as illustrated in Fig. 4. First, the panchromatic image is low-pass filtered by band-specific MTF filters, then down- and up-sampled. Subtracting this result from the original panchromatic image yields the spatial details for injection. The injection coefficients, which control how much detail is added to each multispectral band, are estimated by a regression procedure.
We embed the main components of MTF-GLP into the neural network architecture. The per-band low-pass filtering is implemented as fixed convolutional filters mimicking the sensor’s MTF, corresponding to the panchromatic degradation step in the classical model. GLP-style detail extraction is realized by subtracting the low-pass filtered panchromatic image from the original panchromatic image, with a residual detail network learning only the correction beyond the classical detail term. The regression-based injection mechanism is a gain estimator, initialized around unity, which adapts the detail injection for each band. Finally, a shallow refinement network enforces inter-band consistency, compensating for nonlinear effects not captured by linear regression.
@@ III-B1 Training Strategy and Degradation Simulation
The fusion model is trained in a fully self-supervised manner using only native Sentinel-2 data. We adopt a Wald protocol training strategy where the input bands are downsampled from their native resolution, and the original band serves as the supervision target.
To simulate realistic downsampled images during training, each input band undergoes a degradation process
$\mathbf{X}_{\text{LR}}^{\downarrow}=\left[\Phi_{\boldsymbol{\sigma}}\left(\mathbf{X}_{\text{LR}}\right)\right]_{\downarrow},$ (12)
where $\Phi_{\boldsymbol{\sigma}}(\cdot)$ is an optional isotropic Gaussian blur is applied to emulate sensor response, and $\downarrow$ denotes the boxcar filter, implemented as average pooling with stride corresponding to the target scale.
@@ III-B2 Network Implementation
The fusion model $F_{\boldsymbol{\nu}}(\cdot,\cdot)$ is instantiated via a separate network for each resolution group (10 m, 20 m, and 60 m), with approximately 200k parameters per model.
The model is trained to minimize an $\ell_{1}$ reconstruction loss
$\mathcal{L}_{\text{fusion}}=\left\|\widehat{\mathbf{X}}_{\text{LR}}-\mathbf{X}_{\text{LR}}\right\|_{1},$
where
@@
$\widehat{\mathbf{X}}_{\text{LR}}=\mathcal{F}_{\boldsymbol{\nu}}\left(\mathbf{X}_{\text{LR}}^{\downarrow},\mathbf{x}_{\text{LR}}\right)$ (13)
is the predicted fused output.
Detailed architecture
The details of architecure given in Fig. 4) are as follows:
(1) Convex mixer: The bands of the Sentinel-2 RGB, $\mathbf{x}_{LR}$, is combined into a panchromatic proxy,
$\mathbf{p}=\sum_{c=1}^{3}w_{c}\,\mathbf{x}_{\mathrm{LR}},\qquad w_{c}\geq 0,\ \sum_{c}w_{c}=1,$
where $\{w_{c}\}$ are softmax-constrained, learnable convex weights.
(2) Sensor degradation model: For each band $i$, a Gaussian MTF blur $\Phi_{\sigma_{i}}$ with reflection padding, followed by area downsampling and bicubic upsampling ($\mathcal{D}_{r_{i}},\mathcal{U}_{r_{i}}$), produces the low-pass panchromatic image approximation
$\widetilde{\mathbf{p}}_{i}=\mathcal{U}_{r_{i}}\!\left(\mathcal{D}_{r_{i}}\left(\left(\Phi_{\sigma_{i}}(\mathbf{p}\right)\right)\right),$
and the high frequency details that need to be injected in the multi-spectral bands for each band is given as
$\boldsymbol{\Delta}_{i}=\mathbf{p}-\widetilde{\mathbf{p}}_{i}.$
Here, the scaling factor $r_{i}$ denotes the ratio between the native ground sampling distance (GSD) of the $i$-th multispectral band and the panchromatic proxy resolution. In practice, $r_{i}$ determines the downsampling and upsampling ratio applied in the simulated sensor degradation model (Fig. 4). Specifically,
$r_{i}=\frac{\text{native GSD of band }i}{\text{target PAN GSD}}.$
For the Sentinel-2 configuration, we use three dedicated fusion branches corresponding to the 10 m, 20 m, and 60 m resolution groups. We use $r_{i}=4,8,\text{ and }24$ for the 10 m, 20 m, and 60 m bands, respectively. Each branch reconstructs its respective multispectral group to the 2.5 m target resolution, guided by the super-resolved RGB proxy. Each blur kernel $\Phi_{\sigma_{i}}$ is initialized from the nominal MTF value of the corresponding Sentinel-2 band, converted to a Gaussian standard deviation $\sigma_{i}$ as a function of the Nyquist frequency and the scale ratio $r_{i}$.
(3) DetailNet: Each band’s detail is refined by a small residual network that operates on the concatenation of the GLP detail and the upsampled multispectral band group that is input to the fusion module, $\widetilde{\mathbf{x}}$ (Fig. 4):
[…]
where DetailNet consists of a $3{\times}3$ convolution, LeakyReLU, two RRDB blocks, and a linear $3{\times}3$ conv that outputs a one-channel residual map. The final layer has no activation to retain the signed nature of the learned residual.
(4) GainNet: A per-band residual gain map is estimated from the concatenation of the panchromatic proxy, its band-specific low-pass approximation, and the mean of the upsampled multispectral inputs:
$g_{i}=1+{\rm GainNet}\!\left(\left[\ \mathbf{p},\ \widetilde{\mathbf{p}}_{i},\ \overline{\mathbf{x}}\ \right]\right),$
where $\overline{\mathbf{x}}$ denotes the per-pixel mean of the upsampled multispectral group $\widetilde{\mathbf{x}}$. GainNet consists of a shared two-layer convolutional subnetwork: a $3{\times}3$ convolution with LeakyReLU activation ($0.2$ slope) followed by a $1{\times}1$ convolution that outputs a single-channel residual map $\Delta g_{i}$. The final gain is formulated as $g_{i}=1+\Delta g_{i}$, providing an unbounded residual gain around unity that scales the injected detail adaptively for each band.
(5) RefineNet and fusion: The bandwise fusion applies the gain-modulated detail injection
$\mathbf{y}_{i}=\widetilde{\mathbf{x}}_{i}+g_{i}\cdot\widehat{\mathbf{\Delta}}_{i},$
where $\widetilde{\mathbf{x}}_{i}$ is the $i$th input multispectral band. The output $\{\mathbf{y}_{i}\}$ is then stacked into into $\mathbf{Y}$. A shallow refinement network then enforces inter-band consistency using the concatenation $\mathbf{Y},\ \overline{\boldsymbol{\Delta}}_{i}]$, where $\overline{\boldsymbol{\Delta}}_{i}$ is the mean refined across all detail bands $\{\boldsymbol{\Delta}_{i}\}$:
$\widehat{\mathbf{X}}_{\text{HR}}=\mathbf{Y}+{\rm RefineNet}\!\left(\left[\mathbf{Y},\ \overline{\boldsymbol{\Delta}}_{i}\right]\right),$
RefineNet consists of a $3{\times}3$ convolution with LeakyReLU activation, two residual blocks, and a final linear $3{\times}3$ layer that outputs a residual correction added back to $\mathbf{Y}$.
Inference.
At inference, the fusion model receives inputs at their original resolutions (10 m, 20 m, or 60 m), but the native RGB input is replaced by a 2.5 m GSD super-resolved RGB image. This new input is low-pass filtered with the same per-band MTF kernels used in training to ensure frequency consistency. The network must generalize zero-shot to this enhanced spatial prior, as the super-resolved image is not used during the training phase [lanaras2018super].